# On the Identifiability of Tensor Ranks via Prior Predictive Matching

### v1

@@ Acknowledgements
ES thanks the contributions and discussions with Aaron Schein and Caterina De Bacco that was fundamental for the development of this work. AK belongs to the ELLIS Institute Finland, and acknowledges support from the Research Council of Finland Flagship programme: Finnish Center for Artificial Intelligence FCAI and grant 363317. DM and ES acknowledge the support by the Silicon Valley Community Foundation (SVCF) through the Ripple Impact Fund, the Fundação de Amparo à Pesquisa do Estado do Rio de Janeiro (FAPERJ) through the Jovem Cientista do Nosso Estado program, and the Fundação de Amparo à Pesquisa do Estado de São Paulo (FAPESP) through the grant 2023/00815-6. IU and ES acknowledge the support of “la Caixa” Foundation’s LCF/BQ/PI22/11910028 award, , as well as funds by MICIU/AEI/10.13039/501100011033 and the BERC 2022-2025 program funded by the Basque Government.
@@ A.1 Preliminaries: observation model, covariances, and pure terms
Let $\bm{Y}\in\mathbb{R}^{N_{1}\times\dots\times N_{M}}$ be an observed tensor with $M$ modes. A mode of an order-$M$ tensor is one of its $M$ dimensions. Let $\mathbf{i}=(i_{m})_{m\in[M]}\in\bigtimes_{m=1}^{M}[N_{m}]$ be the multi-index from the observed tensor domain, where $N_{m}]=\{1,\dots,N_{m}\}$ . Let $\mathbf{r}=(r_{p})_{p\in[R]}$ be the multi-rank vector of size $R\leq M$. Let the complete multi-index be $(i_{1},\dots,i_{R})\in\bigtimes_{p=1}^{R}[r_{p}]$ for the product of all $R$ latent rank domains $r_{p}]$, and let $\mathbf{i}_{\beta}$ be a derived multi-index for latent domains $\beta\in\mathcal{K}\subset[R]$ representing a subset $\mathcal{K}$ of all possible latent dimensions. For multi-indices $\mathbf{i},\mathbf{j}$, the shared-mode set is $S(\mathbf{i},\mathbf{j})=\{k:i_{k}=j_{k}\}$. We use bold sans-serif for tensors ($\bm{\eta}$) and vectors/matrices ($\bm{\theta}^{(p)}$). All latent factor multi-indices $\mathbf{i}_{p}(\mathbf{i},\beta)$ are formed by combining some components of the observed tensor indices $\mathbf{i}$ and some components latent domains $\mathbf{i_{\beta}}$, and each factor is probabilistically defined as $\theta^{(p)}_{\mathbf{i}_{p}(\mathbf{i},\beta)}\sim\pi_{p}(\mu_{p},\sigma_{p}^{2})$, i.i.d. across indices with $p\mathcal{P}$ groups the factors/cores. The notation $\mathbf{i}_{p}(\mathbf{i},\beta)$ is small change (a explicit dependence on observation tensor multi-index) on the main article notation $\mathbf{i}_{p}(\alpha,\beta)$ (section II) for convenience in some calculation in the Appendix.
@@ A.1.1 A counting lemma for rank exponents
Expand $\mathbb{E}[\eta_{\mathbf{i}}\eta_{\mathbf{j}}]$ into a double sum over latent indices for the two entries. Independence across different latent tuples makes the expectation factor across parameter groups $p$. Whether the same random variable appears twice (giving $\sigma_{p}^{2}$) or two independent copies (giving $\mu_{p}^{2}$) is determined by index equalities induced by $S$ and by whether the two latent index tuples coincide on the arguments of $\theta^{(p)}$. The number of independent sums that remain for each latent link determines the power of the corresponding rank. Subtracting lower-order $C_{T}$ by inclusion–exclusion isolates the pure term $v_{S}$, which retains a single monomial. ∎
@@ A.2 Theoretical analysis of the Tucker model
which is the clarified version of Eq. (3.1) in the main text (Lemma III.2/Lemma III.3 there).
@@ A.3.2 Detailed TT moment calculations (interior and boundary cases)
Interior patterns (full monomials)
For 1<p<M11<p<M-1:
@@
Worked example (rank exponents table)
For illustration, consider $M=6$ and $p=3$. The link exponents $\alpha_{\ell}(S)$ from Proposition A.4 are:
@@ A.4.1 Model and notation
A.4.2 TR prior predictive moments and rank–ratio identity
Throughout this section, indices are taken modulo $M$. Each TR core $\Theta^{(p)}_{i_{p}}\in\mathbb{R}^{r_{p-1}\times r_{p}}$ has i.i.d. entries with
$\mathbb{E}[\theta^{(p)}_{ab}]=\mu_{p},\qquad\mathrm{Cov}(\theta^{(p)}_{ab},\theta^{(p)}_{a^{\prime}b^{\prime}})=\sigma_{p}^{2}\,\delta_{aa^{\prime}}\delta_{bb^{\prime}}.$
Conditional on $\bm{\eta}$, the observations satisfy $\mathbb{E}[Y_{\mathbf{i}}\mid\bm{\eta}]=\eta_{\mathbf{i}}$.
Lemma A.9 (TR prior predictive moments under exact sharing).
For any $M\geq 3$,
@@ A.4.2 TR prior predictive moments and rank–ratio identity
Mean. Expanding the trace, $\mathbb{E}[\eta_{\mathbf{i}}]=\sum_{s_{0},\dots,s_{M-1}}\prod_{p=1}^{M}\mathbb{E}[\theta^{(p)}_{s_{p-1},i_{p},s_{p}}]=\prod_{p}(r_{p}\mu_{p})$, giving equation 29.
Singleton $v_{\{p\}}$. Consider $\mathrm{Cov}(\eta_{\mathbf{i}},\eta_{\mathbf{j}})$ for the exact-sharing pattern $S=\{p\}$ (only mode $p$ shared, all others free). In $\mathbb{E}[\eta_{\mathbf{i}}\eta_{\mathbf{j}}]$, modes $k\neq p$ contribute $\mu_{k}^{2}$, while mode $p$ contributes $\sigma_{p}^{2}$. The rank exponents count independent latent sums: links $k\notin\{p-1,p\}$ are free ($r_{k}^{2}$); the incoming link p1p-1 is jointly summed ($r_{p-1}$); the outgoing link $p$ remains decoupled ($r_{p}^{2}$). This yields equation 30.
Adjacent pair $v_{\{p,p+1\}}$. For the exact-sharing pattern $S=\{p,p+1\}$, only cores $p$ and $p+1$ are in variance, others in mean. Two consecutive variances glue the two copies into a single cyclic loop: all latent links participate in exactly one joint sum, hence each $r_{k}$ appears once. Multiplying by $\sigma_{p}^{2}\sigma_{p+1}^{2}\prod_{k\notin\{p,p+1\}}\mu_{k}^{2}$ gives equation 31. ∎
Remarks.
(i) Mean normalization of pure terms leaves $\xi_{p}$ unchanged but improves numerical stability. (ii) Edge cases: $M=2$ reduces to a CP/matrix case; $M=3$ has eigenvalues $1,1\pm i\sqrt{3}$ and remains identifiable.
@@
$\boxed{\qquad\xi_{p}\;=\;r_{p}\,\frac{r_{p-1}}{r_{p+1}},\qquad}$ (32)
defining a circulant system $\mathbf{C}\mathbf{x}=\bm{\psi}$ for $x_{p}=\log r_{p}$ and $\psi_{p}=\log\xi_{p}$.
@@
Substituting equation 29–equation 31, all $\mu_{k}$ and $\sigma_{k}$ cancel, leaving $\xi_{p}=r_{p}\,\frac{r_{p-1}}{r_{p+1}}$. Taking logs with $x_{p}=\log r_{p}$ gives the circulant system
$x_{p}+x_{p-1}-x_{p+1}=\psi_{p},\quad p=1,\ldots,M.$ (33)
@@
The coefficient matrix of equation 33 is circulant with first row $(1,-1,0,\dots,0,1)$. Its DFT eigenvalues are
$\lambda_{k}=1+2i\sin(2\pi k/M),\quad k=0,\dots,M-1,$
all nonzero for $M\geq 3$. Hence $\mathbf{C}$ is invertible and the ranks are uniquely recovered by $\mathbf{x}=\mathbf{C}^{-1}\bm{\psi}$, $r_{p}=\exp(x_{p})$. ∎
@@
Remarks.
(i) Mean normalization of pure terms leaves $\xi_{p}$ unchanged but improves numerical stability. (ii) Edge cases: $M=2$ reduces to a CP/matrix case; $M=3$ has eigenvalues $1,1\pm i\sqrt{3}$ and remains identifiable.
@@ B.1 Covariances, interaction sets, and pure terms
We estimate ranks in two steps. (i) Observable covariances. For a sharing set $S\subseteq\{1,\dots,M\}$, the exact–sharing covariance $\widehat{C}_{S}$ is computed from pairs $(Y_{\mathbf{i}},Y_{\mathbf{j}})$ with indices equal on $S$ and different on the complement. (ii) Pure terms. The pure interaction $\widehat{v}_{S}$ is obtained from $\{\widehat{C}_{T}:T\subseteq S\}$ via Möbius inversion:
$\widehat{v}_{S}\;=\;\sum_{T\subseteq S}(-1)^{|S|-|T|}\,\widehat{C}_{T}.$
@@
Complexity Analysis (Simplified). Let $\mathcal{S}$ be the number of distinct sharing sets used, $B$ bootstraps, and $S_{\text{cov}}$ pairs per set. Pair generation and access are $O(1)$ amortized; Möbius is $O(1)$ per set since $|S|\leq 3$ here. The dominant cost is sampling:
@@ B.2 Variance and SNR of the Pure-Term Estimators
Let $v_{S}$ be the TT pure interaction defined by exact–sharing Moebius inversion over covariances $C_{T}$. With per-mode prior mean $\mu_{m}$, variance $\sigma_{m}^{2}$, second raw moment $m_{2,m}=\mu_{m}^{2}+\sigma_{m}^{2}$ and fourth raw moment $m_{4,m}$, the TT population pure term is
@@
Let $\widehat{C}_{T}$ be the exact–sharing covariance estimated from $n$ paired samples and $\widehat{v}_{S}=\sum_{T\subseteq S}(-1)^{|S|-|T|}\widehat{C}_{T}$. Under independent paired sampling per TT and large $n$,
$\mathrm{Var}(\widehat{C}_{T})\approx\frac{1}{n}\Bigg(\prod_{\ell}r_{\ell}^{\,h_{\ell}(T)}\Bigg)\Bigg(\prod_{m\in T}m_{4,m}\Bigg)\Bigg(\prod_{m\notin T}m_{2,m}^{2}\Bigg)-\frac{1}{n}\Bigg(\prod_{\ell}r_{\ell}^{\,2e_{\ell}(T)}\Bigg)\Bigg(\prod_{m\in T}m_{2,m}^{2}\Bigg)\Bigg(\prod_{m\notin T}\mu_{m}^{4}\Bigg),$
where $h_{\ell}(T)=2e_{\ell}(T)\in\{2,4\}$ are the fourth–moment exponents for TT. Consequently,
$\mathrm{Var}(\widehat{v}_{S})\;\approx\;\sum_{T\subseteq S}\frac{1}{n}(-1)^{2|S|-2|T|}\left[\Bigg(\prod_{\ell}r_{\ell}^{\,h_{\ell}(T)}\Bigg)\Bigg(\prod_{m\in T}m_{4,m}\Bigg)\Bigg(\prod_{m\notin T}m_{2,m}^{2}\Bigg)-\Bigg(\prod_{\ell}r_{\ell}^{\,2e_{\ell}(T)}\Bigg)\Bigg(\prod_{m\in T}m_{2,m}^{2}\Bigg)\Bigg(\prod_{m\notin T}\mu_{m}^{4}\Bigg)\right].$
@@
We can observe from Figure 2 that the SNR of the pure-interaction covariance estimator for $|S|=1$ and $|S|=2$ (different sizes of interaction sets) has a narrow band where it lead to joint high SNR. In most areas of the graph we have prior configurations that leads to low SNR, affecting the quality and variability of any estimator based on pure-interactin covariance values.

### v2

@@ Acknowledgements
We would like to thank Caterina de Bacco, Aaron Schein, and Tomasz Kuśmierczyk for their insightful dialog and constructive feedback provided during the early stages of this research.
E. da Silva. and I. Urteaga were supported by “la Caixa” foundation’s LCF/BQ/PI22/11910028 award, as well as the Basque Government through the BERC 2022-2025 program and the Ministry of Science and Innovation’s BCAM Severo Ochoa accreditation CEX2021-001142-S/MICIN/AEI/10.13039/501100011033.
D. Mesquita acknowledges the support by the Fundação Carlos Chagas Filho de Amparo à Pesquisa do Estado do Rio de Janeiro (FAPERJ) (SEI-260003/020348/2025, SEI-260003/020694/2025) and the Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) (404336/2023-0, 305692/2025-9).
E. da Silva’s work is funded by Portuguese national funds through FCT – Foundation for Science and Technology, I.P., within the scope of the research unit UID/00326 - Centre for Informatics and Systems of the University of Coimbra.
I. Urteaga also acknowledges the support of Grant RYC2023-045922-I funded by MICIU/AEI/10.13039/501100011033 and by ESF+.
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes]
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [Yes]
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. [Yes]
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. [Yes]
    2. (b)
      Complete proofs of all theoretical results. [Yes] We include some proofs in the main text and other in the appendix.
    3. (c)
      Clear explanations of any assumptions. [Yes]
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). [Yes]
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). [Yes]
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). [Yes]
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). [Yes]
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. [Not Applicable]
    2. (b)
      The license information of the assets, if applicable. [Not Applicable]
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. [Not Applicable]
    4. (d)
      Information about consent from data providers/curators. [Not Applicable]
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. [Not Applicable]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. [Not Applicable]
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. [Not Applicable]
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. [Not Applicable]
On the Identifiability of Tensor Ranks via Prior Predictive Matching:
Supplementary Materials
@@ A.1 Preliminaries: observation model, covariances, and pure terms
Let $\bm{Y}\in\mathbb{R}^{N_{1}\times\dots\times N_{M}}$ be an observed tensor with $M$ modes. A mode of an order-$M$ tensor is one of its $M$ dimensions. Let $\mathbf{i}=(i_{m})_{m\in[M]}\in\bigtimes_{m=1}^{M}[N_{m}]$ be the multi-index from the observed tensor domain, where $N_{m}]=\{1,\dots,N_{m}\}$ . Let $\mathbf{r}=(r_{p})_{p\in[R]}$ be the multi-rank vector of size $R\leq M$. Let the complete multi-index be $(i_{1},\dots,i_{R})\in\bigtimes_{p=1}^{R}[r_{p}]$ for the product of all $R$ latent rank domains $r_{p}]$, and let $\mathbf{i}_{\beta}$ be a derived multi-index for latent domains $\beta\in\mathcal{K}\subset[R]$ representing a subset $\mathcal{K}$ of all possible latent dimensions. For multi-indices $\mathbf{i},\mathbf{j}$, the shared-mode set is $S(\mathbf{i},\mathbf{j})=\{k:i_{k}=j_{k}\}$. We use bold sans-serif for tensors ($\bm{\eta}$) and vectors/matrices ($\bm{\theta}^{(p)}$). All latent factor multi-indices $\mathbf{i}_{p}(\mathbf{i},\beta)$ are formed by combining some components of the observed tensor indices $\mathbf{i}$ and some components latent domains $\mathbf{i_{\beta}}$, and each factor is probabilistically defined as $\theta^{(p)}_{\mathbf{i}_{p}(\mathbf{i},\beta)}\sim\pi_{p}(\mu_{p},\sigma_{p}^{2})$, i.i.d. across indices with $p\in\mathcal{P}$ groups the factors/cores.
@@ A.1.1 A counting lemma for rank exponents
Expand $\mathbb{E}[\eta_{\mathbf{i}}\eta_{\mathbf{j}}]$ into a double sum over latent indices for the two entries. Independence across different latent tuples makes the expectation factor across parameter groups $p$. Whether the same random variable appears twice (giving $\sigma_{p}^{2}$) or two independent copies (giving $\mu_{p}^{2}$) is determined by index equalities induced by $S$ and by whether the two latent index tuples coincide on the arguments of $\theta^{(p)}$. The number of independent sums that remain for each latent link determines the power of the corresponding rank. Subtracting lower-order $C_{S^{\prime}}$ by inclusion–exclusion isolates the pure term $v_{S}$, which retains a single monomial. ∎
@@ A.2 Theoretical analysis of the Tucker model
which is the clarified version of Eq. (3.1) in the main text (Lemma 3.2/Lemma 3.3 there).
@@ A.3.2 Detailed TT moment calculations (interior and boundary cases)
Interior patterns (full monomials) For 1<p<M11<p<M-1:
@@
Worked example (rank exponents table) For illustration, consider $M=6$ and $p=3$. The link exponents $\alpha_{\ell}(S)$ from Proposition A.4 are:
@@ A.4.1 Model and notation
A.4.2 TR moments: monomials and rank exponents
We first obtain the three families of monomials used throughout: $(E[Y])^{2}$, singletons $v_{\{p\}}$, and adjacent pairs $v_{\{p,p+1\}}$.
Lemma A.9 (TR prior predictive moments).
For any $M\geq 3$ (indices modulo $M$), the following hold:
@@ A.4.2 TR moments: monomials and rank exponents
Mean. From equation 24, $\mathbb{E}[\eta_{\mathbf{i}}]=\sum_{s}\prod_{p}\mu_{p}=\prod_{p}(r_{p}\mu_{p})$, hence equation 25.
Singleton $v_{\{p\}}$. Consider $\mathrm{Cov}(\eta_{\mathbf{i}},\eta_{\mathbf{j}})$ for pairs sharing all modes and isolate the pure contribution where only $\Theta^{(p)}$ is in variance (Appendix A.1). Expanding $\mathbb{E}[\eta_{\mathbf{i}}\eta_{\mathbf{j}}]$ yields a double sum over latent chains $(s_{0},\dots,s_{M-1})$ and $(t_{0},\dots,t_{M-1})$. Modes $k\neq p$ contribute $\mu_{k}^{2}$; mode $p$ contributes $\sigma_{p}^{2}$. The rank exponents count independent sums per link:
  • Links not adjacent to $p$ (i.e., $k\notin\{p-1,p\}$) are unconstrained in both copies $\Rightarrow$ 2 independent sums $\Rightarrow r_{k}^{2}$.
  • The link p1p-1 is adjacent to the sole variance and becomes constrained across copies by equality of the index that flows into $\Theta^{(p)}$; only one joint sum remains $\Rightarrow r_{p-1}$.
  • The link $p$ (outgoing from $\Theta^{(p)}$) remains unconstrained across copies because the variance identification at $p$ does not couple the outgoing indices in the two copies; this keeps two independent sums $\Rightarrow r_{p}^{2}$.
This gives equation 26.
Adjacent pair $v_{\{p,p+1\}}$. Now only $\Theta^{(p)}$ and $\Theta^{(p+1)}$ are in variance, all other modes in mean. In the ring, placing two consecutive variances glues the two copies into a single cyclic loop: every link becomes constrained across the copies, leaving exactly one independent sum per link. Hence each $r_{k}$ appears to the first power, and the moment factor is $\sigma_{p}^{2}\sigma_{p+1}^{2}\prod_{k\notin\{p,p+1\}}\mu_{k}^{2}$, yielding equation 27. ∎
Discussion.
The asymmetric exponents in equation 26 (power 1 on $r_{p-1}$, but power 2 on $r_{p}$) are a genuine ring effect: the variance at node $p$ constrains the incoming link (shared across copies) but leaves the outgoing link with independent sums. In contrast, two adjacent variances collapse the two-copy ring globally, forcing single powers for all links in equation 27.
A.4.3 Rank–ratio identity and identifiability
Define the observable statistic
$\xi_{p}\;:=\;\frac{(E[Y])^{2}\;v_{\{p,p+1\}}}{v_{\{p\}}\;v_{\{p+1\}}}.$ (28)
@@ A.4.3 Rank–ratio identity and identifiability
$\boxed{\qquad\xi_{p}\;=\;r_{p}\,\frac{r_{p-1}}{r_{p+1}}.\qquad}$ (29)
@@
Substitute equation 25–equation 27 into equation 28. All $\mu$ and $\sigma$ factors cancel. For ranks, the numerator contributes $\big(\prod_{k}r_{k}\big)\cdot\big(\prod_{k}r_{k}\big)$, while the denominator contributes $(\prod_{k\notin\{p-1,p\}}r_{k}^{2})r_{p-1}r_{p}^{2}\cdot(\prod_{k\notin\{p,p+1\}}r_{k}^{2})r_{p}r_{p+1}^{2}$. After cancellation, only the factor $r_{p}r_{p-1}/r_{p+1}$ remains. ∎
Let $x_{p}=\log r_{p}$ and $\psi_{p}=\log\xi_{p}$. Taking logs in equation 29 gives the circulant system
$x_{p}+x_{p-1}-x_{p+1}\;=\;\psi_{p},\qquad p=1,\dots,M,$ (30)
@@
Let $\omega=e^{2\pi i/M}$. The eigenpairs of a circulant are obtained by the DFT: the $k$th eigenvalue is
$\lambda_{k}\;=\;1+\omega^{k}-\omega^{-k}\;=\;1+2i\,\sin(2\pi k/M),\qquad k=0,\dots,M-1.$
For $M\geq 3$, $\sin(2\pi k/M)=0$ iff $k\in\{0,M\}$; in either case $\lambda_{k}=1\neq 0$. Thus no $\lambda_{k}$ vanishes and $\det\mathbf{C}\neq 0$. Therefore equation 30 has a unique solution $\mathbf{x}$. Exponentiating gives the ranks. ∎
@@
A.4.4 Detailed TR moment calculations and edge cases
We briefly expand the counting in Lemma A.9.
Singleton $v_{\{p\}}$.
Write the two copies of the ring with latent chains $(s_{0},\dots,s_{M})$ and $(t_{0},\dots,t_{M})$ (with $s_{0}=s_{M}$, $t_{0}=t_{M}$). The variance at $p$ identifies the same entry of $\Theta^{(p)}$ across copies, which forces $s_{p-1}=t_{p-1}$ but leaves $s_{p}$ and $t_{p}$ independent. Propagating around the cycle:
  • For links far from $p$ (neither endpoint adjacent to $p$), both copies retain independent sums $\Rightarrow r_{k}^{2}$.
  • At link p1p-1 the identification $s_{p-1}=t_{p-1}$ leaves one sum $\Rightarrow r_{p-1}$.
  • At link $p$ the two outgoing indices remain independent across copies $\Rightarrow r_{p}^{2}$.
Moment factors follow since only mode $p$ contributes $\sigma_{p}^{2}$; all others contribute $\mu_{k}^{2}$. This reproduces equation 26.
Adjacent pair $v_{\{p,p+1\}}$.
Variance at $p$ and $p+1$ identifies $\Theta^{(p)}$ and $\Theta^{(p+1)}$ across copies. In the ring, this creates a single joint cycle: starting from any link, equality constraints propagate all the way around, leaving exactly one independent sum per link $\Rightarrow\prod_{k}r_{k}$. The moment factor is $\sigma_{p}^{2}\sigma_{p+1}^{2}\prod_{k\notin\{p,p+1\}}\mu_{k}^{2}$, giving equation 27.
Small-$M$ edge cases.
  • $M=2$: the TR degenerates into a matrix factorization with a single rank parameter; the CP formulas apply.
  • $M=3$: all formulas above hold; $\mathbf{C}$ has eigenvalues $1,\,1\pm i\sqrt{3}$ (all nonzero), so identifiability still holds.
Robustness to minor model variants.
If the entries of $\Theta^{(p)}$ are centered ($\mu_{p}=0$), the above uses $(E[Y])^{2}$ only via $\prod\mu_{k}$, which would be zero; in that case one can replace $(E[Y])^{2}$ with $\mathrm{Var}(\eta)$ and form analogous ratios using higher moments. Our main identifiability result relies on $\mu_{p}\neq 0$, as assumed in the main text.
@@ B.1 Covariances, interaction sets, and pure terms
We estimate ranks in two steps.
  1. 1.
    Observable covariances. For a sharing set $S\subseteq\{1,\dots,M\}$, the exact–sharing covariance $\widehat{C}_{S}$ is computed from pairs $(Y_{\mathbf{i}},Y_{\mathbf{j}})$ with indices equal on $S$ and different on the complement.
  2. 2.
    Pure terms. The pure interaction $\widehat{v}_{S}$ is obtained from $\{\widehat{C}_{S^{\prime}}:{S^{\prime}}\subseteq S\}$ via Möbius inversion:
    $\widehat{v}_{S}\;=\;\sum_{{S^{\prime}}\subseteq S}(-1)^{|S|-|{S^{\prime}}|}\,\widehat{C}_{S^{\prime}}.$
@@
B.1.1 Complexity Analysis (Simplified).
Let $\mathcal{S}$ be the number of distinct sharing sets used, $B$ bootstraps, and $S_{\text{cov}}$ pairs per set. Pair generation and access are $O(1)$ amortized; Möbius is $O(1)$ per set since $|S|\leq 3$ here. The dominant cost is sampling:
@@ B.2 Variance and SNR of the Pure-Term Estimators
Let $v_{S}$ be the TT pure interaction defined by exact–sharing Moebius inversion over covariances $C_{S}$. With per-mode prior mean $\mu_{m}$, variance $\sigma_{m}^{2}$, second raw moment $m_{2,m}=\mu_{m}^{2}+\sigma_{m}^{2}$ and fourth raw moment $m_{4,m}$, the TT population pure term is
@@
Let $\widehat{C}_{S^{\prime}}$ be the exact–sharing covariance estimated from $n$ paired samples and $\widehat{v}_{S}=\sum_{{S^{\prime}}\subseteq S}(-1)^{|{S^{\prime}}|-|{S^{\prime}}|}\widehat{C}_{S^{\prime}}$. Under independent paired sampling per $S^{\prime}$} and large $n$,
$\mathrm{Var}(\widehat{C}_{S^{\prime}})\approx\frac{1}{n}\Bigg(\prod_{\ell}r_{\ell}^{\,h_{\ell}({S^{\prime}})}\Bigg)\Bigg(\prod_{m\in{S^{\prime}}}m_{4,m}\Bigg)\Bigg(\prod_{m\notin{S^{\prime}}}m_{2,m}^{2}\Bigg)-\frac{1}{n}\Bigg(\prod_{\ell}r_{\ell}^{\,2e_{\ell}({S^{\prime}})}\Bigg)\Bigg(\prod_{m\in{S^{\prime}}}m_{2,m}^{2}\Bigg)\Bigg(\prod_{m\notin{S^{\prime}}}\mu_{m}^{4}\Bigg),$
where $h_{\ell}({S^{\prime}})=2e_{\ell}(T{S^{\prime}}\in\{2,4\}$ are the fourth–moment exponents for TT. Consequently,
$\mathrm{Var}(\widehat{v}_{S})\;\approx\;\sum_{{S^{\prime}}\subseteq S}\frac{1}{n}(-1)^{2|S|-2|{S^{\prime}}|}\left[\Bigg(\prod_{\ell}r_{\ell}^{\,h_{\ell}({S^{\prime}})}\Bigg)\Bigg(\prod_{m\in{S^{\prime}}}m_{4,m}\Bigg)\Bigg(\prod_{m\notin{S^{\prime}}}m_{2,m}^{2}\Bigg)-\Bigg(\prod_{\ell}r_{\ell}^{\,2e_{\ell}({S^{\prime}})}\Bigg)\Bigg(\prod_{m\in{S^{\prime}}}m_{2,m}^{2}\Bigg)\Bigg(\prod_{m\notin{S^{\prime}}}\mu_{m}^{4}\Bigg)\right].$
@@
We can observe from Figure 2 that the SNR of the pure-interaction covariance estimator for $|S|=1$ and $|S|=2$ (different sizes of interaction sets) has a narrow band leading to joint high SNR. In most areas of the graph we have prior configurations that lead to low SNR, affecting the quality and variability of any estimator based on pure-interactin covariance values.